# Pi in the Sky: Neutron Stars with Exceptionally Light QCD Axions

### v1

@@ 2.2 Effective axion potential at finite density
For $\varepsilon=1$, the critical density determined from Eq. (11) is sufficiently large that the expansion in $\sigma_{N}n_{B}/m_{\pi}^{2}f_{\pi}^{2}$ breaks down near the critical density, and it is conceivable that axion condensation does not occur at any density. We will return to this question in Sec. 3.2. However, for $\varepsilon<0.4$, Eq. (10) predicts axion condensation at $n_{B}\lesssim n_{\rm sat}$. At these moderate densities where matter can be described reliably using nuclear physics, one can conclude that the axion potential is minimized at $\theta=\pi$. If the nuclear force is significantly modified by axion condensation, $\theta=\pi$ may be favored below the saturation density even for $\varepsilon\geq 0.4$, which we will discuss in more detail in Sec. 3 and Sec. 4.2.
@@
where $P(\mu_{B},\theta)$ is the pressure, $k_{F}(\theta)=\sqrt{\mu_{B}^{2}-m^{2}_{n}(\theta)}$ and the neutron mass $m_{n}(\theta)$ was defined earlier in Eq. (2.1). From Eq. (12) we see that axion contribution to the pressure is negative, and for $\theta=\pi$, $P_{\rm axion}=-V(\theta=\pi)=\varepsilon f_{\pi}^{2}[m_{\pi}^{2}(\theta=\pi)-m_{\pi}^{2}(\theta=0)]$. The importance of this negative pressure was first pointed out in Ref. Csaki:2018fls ; Balkin:2022qer where it was noted that for white dwarfs, which are stabilized by electron degeneracy pressure, this negative pressure would destabilize a large number of observed white dwarfs in the universe. This consideration leads to the strongest constraint on the axion parameter space for $\varepsilon<2\times 10^{-7}$ Balkin:2022qer . A neutron star, on the other hand, is much denser and balanced by the much larger neutron degeneracy pressure. It is unlikely that this negative pressure can lead to the collapse of a neutron star into black holes, but it might significantly change the EOS of a neutron star.
@@
We begin by analyzing the effects of axion on a dilute gas of neutrons. The left panel of Fig. 1 shows the pressure of a free Fermi gas of neutrons at $\theta=\pi$ and $\mu_{B}=m_{n}(\theta=0)$. The insets show the $\theta$-dependence of the free energy in each of the regions depicted. When the pressure is positive, axion condensation will be favored for dilute neutron matter. In the absence of nuclear interactions, dilute neutron matter with $\theta=0$ is metastable for $0.005\lesssim\varepsilon\lesssim 0.065$ and unstable when $\varepsilon\lesssim 0.005$. For $\varepsilon\gtrsim 0.065$, the first-order transition to the $\theta=\pi$ phase occurs at $\mu_{B}>m_{n}$ and low-density neutron matter with $\theta=0$ is stable. In this case, the phase transition to axion condensed matter will occur at finite density. On the other hand, if $\varepsilon\lesssim 0.065$, the phase transition will occur at $\mu_{B}<m_{n}$, and bulk matter in its ground state will always have $\theta>\theta_{c}$, with $m_{n}(\theta_{c})=\mu_{B}$. In the case where $2\times 10^{-7}<\varepsilon\lesssim 0.005$, normal matter is unstable and ordinary stars, planets, and white dwarfs are stabilized by Coulomb forces between atoms and nuclei. For $\varepsilon>2\times 10^{-7}$, a system of atoms or nuclei at fixed chemical potential must significantly increase its density in order to nucleate the $\theta=\pi$ phase and an electrical barrier must be overcome to bring the nuclei closer together. For smaller $\varepsilon$, the required increase in density is more modest, and the constraint on axions from the stability of white dwarfs, stars, and planets applies.
For a given $\varepsilon$, if the $\mu_{B}$ inside the star interior exceeds the critical chemical potential and it is possible to nucleate the $\theta=\pi$ phase, axions would condense in the inner region and a domain wall with thickness $\delta R\approx 1/m_{a}$ would separate it from the $\theta=0$ phase in the exterior. In the Fermi gas model, for $\varepsilon\gtrsim 0.065$, the neutron star would consist of three regions: a normal outer region with $\theta=0$, a metastable region which could, in principle, exist in the $\theta=0$ phase if the nucleation timescale was long, and inner core characterized by $\theta=\pi$. In the right panel of Fig. 1, we plot pressure as a function of $\mu_{B}$ for $\varepsilon=0.3$. It consists of the three regions just mentioned, and the insets show the $\theta$ dependence of the free energy, which delineates the stable, metastable, and unstable regions.
2.3 Axion field profile interpolating between $\theta=\pi$ and $\theta=0$
@@
where $\mu^{c}_{B}$ is the critical baryon chemical potential at which the transition occurs. This equation, though used here for a dilute gas of nucleons, is exact when nuclear interaction is included in $\Omega(\theta,\mu^{c}_{B})$ in forthcoming sections. Such an axion profile will develop as long as the object that contains this dilute gas of nucleons is both dense enough (see Eq. (11)) and large enough such that the bulk of the star can be subsumed within a domain wall of thickness
$\Delta r_{\rm DW}\simeq\frac{1}{|m_{a}(n_{B})|}\leq r_{\rm NS}.$ (14)
@@
We will focus on the condition on the density in Eq. (11) for now and return to the effects of the finite size of stellar objects in Sec. 5. For axions with $m_{a}<10^{-11}\,{\rm eV}$, the axion profile can be treated as situated outside the neutron star. In binary systems, axion-mediated forces and axion radiation are consequences of this field profile and lead to the most stringent constraint from aLIGO observations of neutron star binaries Zhang:2021mks . On the other hand, when $m_{a}\gg 10^{-11}\,{\rm eV}$, the axion field profile can reside inside the neutron star, separating regions with different density and pressure. Given the prior constraints on the exceptionally light QCD axion parameter space, we are mostly interested in the region of parameter space where $|m_{a}^{2}(n_{B})|>m_{a}^{2}>1/r_{\rm NS}^{2}$.
@@
In this case, the thickness of the region where the axion field changes from ±$\pm\pi f_{a}$ to 0 is much smaller than the radius of a neutron star, while on the other hand, much larger than the size of a nucleon, typical separation between nucleons, Debye screening length, and mean free path of neutrons and protons in a neutron star. As a result, we can treat the axion field profile as smooth when considering interactions between the nucleons, while at the same time, studying the properties of the gas of nucleons in domains where $\theta=\pm\pi$ and $\theta=0$ separately, and glue them together with a boundary condition at the location of a thin axion “domain wall.”
@@
Unlike a normal neutron star where density and pressure are continuous, with an axion field profile, both the density and pressure are no longer continuous when adopting the approximation that the domain wall is infinitely thin. In addition to a negative pressure from the condensation of the axion field itself, the domain wall also exerts a force inward due to domain wall tension of order $m_{a}f_{a}^{2}/r_{\rm NS}$. For $m_{a}\gg 1/r_{\rm NS}$, this additional pressure is parametrically smaller in the thin wall limit in the model in Eq. (5). We return to discuss the importance of domain wall pressure in the $\mathbb{Z}_{\mathcal{N}}$ model in Sec. 4.2.
3 Neutron and nuclear matter at $\theta=\pi$
@@
In this section, we describe our efforts to model the properties of dense nuclear matter at $\theta=\pm\pi$, taking into account changes to the strength of nuclear interactions. Despite developments in Effective Field Theory (EFT) methods to describe nuclear interactions, the connection between fundamental parameters that appear in the QCD Lagrangian (the quark mass matrix ${\cal M}_{q}$ and $\theta$) and the parameters needed to describe interactions between nucleons remains elusive. This is because the quark mass dependence of the low-energy constants (LECs) in the EFT that describe the strength of interactions at short distances are poorly known. Nevertheless, earlier studies have provided useful insights to address the question of the quark mass dependence of the nucleon-nucleon interaction.
@@
We first briefly summarize earlier work in Refs. Beane:2002vs ; Epelbaum:2002gb ; Berengut:2013nh which are based on nuclear potentials derived from ChiEFT as they are better suited to address how nuclear interactions are altered by changes to the quark mass matrix. To assess the quark mass dependence of the nuclear interactions at momentum scales of relevance to large nuclei and neutron-rich matter encountered in neutron stars, we next compare results from ChiEFT and three scenarios in RMFT.
In ChiEFT nuclear forces are derived by explicitly including the contribution from pion-exchanges and the short-distances components are included through a systematic expansion of operators of increasing dimension constructed from nucleon fields consistent with the symmetries of low-energy QCD. Such a low energy EFT description is expected to be applicable for $n_{B}\lesssim 2n_{\rm sat}$, and is described in Sec 3.2. We employ ChiEFT potentials at order N2LO and perform calculations of the energy per particle in pure neutron matter (PNM) using many-body perturbation theory (MBPT). In this case, the quark mass dependence is taken into account through its modification of the pion mass, nucleon masses, $g_{A}$, and $f_{\pi}$.
@@
RMFT is based on the assumption that scalar and vector interactions provide the dominant contributions to the interaction energy of baryons in large systems. In this phenomenological model, the interaction between nucleons is mediated by the exchange of scalar and vector mesons. The model and the quark mass dependence of the model parameters are described in Sec. 3.3. While ChiEFT is a microscopic theory in which potentials are constrained by nucleon-nucleon scattering data and properties of light nuclei, RMFT is phenomenological model designed to provide a good description of large nuclei. RMFT also provides a convenient extrapolation to high density and permits efficient calculation of heterogeneous crust phases. In RMFT Scenario A, the only modification we make to nuclear physics is to modify the nucleon masses according to Eq. (2.1) since the pion does not explicitly appear in RMFT. In Scenario B, we modify the meson masses with $\theta$ dependence given by calculations in Refs. HanhartPelaezRios2008 ; Pelaez:2010sigmamass ; Berengut:2013nh . In Scenario C, we consider a more extreme modification to the RMFT based on the results around saturation density given by ChiEFT.
@@
3.1 Pion mass dependence of two-nucleon observables
By explicitly accounting for the $m_{\pi}$ dependence of the parameters that describe the nucleon-nucleon (NN) interaction in ChiEFT, one can calculate the quark mass dependence of low-energy NN observables such as the NN scattering phase shifts and the deuteron binding energy. In ChiEFT, at leading order (LO) in the momentum expansion, the NN potential is characterized by long-range one-pion exchange (OPE) and a short-range NN contact interaction. While the $m_{\pi}$ dependence of the parameters associated with the OPE potential is well understood, the $m_{\pi}$ dependence of the NN contact interaction, which is typically called $D_{2}$, is not well determined and the resulting uncertainty in the scattering length and scattering phase shifts have been studied in earlier work Beane:2002vs ; Epelbaum:2002gb .
In Ref. Epelbaum:2002gb , the authors find that when the quark mass is reduced, the binding energy of the deuteron increases and the spin-singlet and spin-triplet scattering lengths decrease. To quantify the quark mass dependence of an observable $\cal O$, the authors define the logarithmic derivative at the physical point
@@ 3.1 Pion mass dependence of two-nucleon observables
since at leading order in ChiPT, $m_{\pi}^{2}\propto m_{q}$. The $m_{q}$ dependence of the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ scattering lengths and the deuteron binding energies are captured by the values of $K_{a1S0}$, $K_{a3S1}$, and $K_{\rm deut}$, and their values obtained in earlier studies are shown in Table 1. Results shown in the first row were obtained using potentials derived from ChiEFT at next-to-leading order (NLO) in Weinberg power counting Epelbaum:2002gb and NLO results obtained in Ref. Beane:2002vs using Kaplan-Savage-Wise power counting Kaplan:1998tg are shown in the second row. In this case, the uncertainty is dominated by the poorly known LEC, $D_{2}$, which was mentioned earlier. Results obtained more recently in Ref. Berengut:2013nh are shown in the third row. In this case, the authors employ strategies to go beyond the NLO approach adopted in Refs. Beane:2002vs ; Epelbaum:2002gb and use information from resonance saturation models to constrain the quark mass dependence of the short-range contact operators and incorporate N2LO corrections to $g_{A}$, $f_{\pi}$ and the nucleon mass.
Table 1: Quark mass dependence of the scattering lengths and the deuteron binding energy.
@@
Epelbaum et al. Epelbaum:2002gb $5\pm 5$ $1.1\pm 0.6$ $-2.8\pm 1.2$
Beane et al. Beane:2002vs $2.4\pm 3.0$ $3.0\pm 3.5$ $-7\pm 6$
Berengut et al. Berengut:2013nh $2.3^{+1.6}_{-1.5}$ $0.32^{+0.17}_{-0.18}$ $-0.86^{+0.45}_{-0.50}$
With decreasing quark mass, these earlier studies found the ${}^{1}S_{0}$ phase shifts were reduced and the ${}^{3}S_{1}$ phase shift were enhanced. Thus, in dilute neutron matter where spin singlet interactions dominate, the interaction energy will be less attractive when s-wave interactions dominate. In dilute nuclear matter with an equal number of protons and neutrons (symmetric nuclear matter), where spin triplet interactions dominate, the s-wave interaction energy is larger and more attractive when the quark mass is reduced. However, when the pion becomes lighter and of longer range, higher partial wave contributions are in general enhanced. The attraction in the ${}^{3}P_{0}$ and ${}^{3}P_{2}$ channels is enhanced and the repulsion in the ${}^{1}P_{1}$ channel is reduced Epelbaum:2002gb . For these reasons, the analysis of the scattering lengths and the deuteron binding energy is insufficient to inform about the behavior of matter at densities of relevance to nuclei and neutron stars where contributions due to higher partial waves and higher-order quark mass-dependent short-distance operators can contribute.
3.2 Axion condensation in Chiral Effective Field Theory
To address how nuclear interactions influence axion condensation, we first make the simple observations that axion condensation is favored when the interaction energy per particle decreases with pion mass. From the discussion in Sec. 2, we infer that in this case condensation would occur at $n_{B}\lesssim 2.6n_{\rm sat}$ for $\varepsilon=1$. In ChiEFT, the $m_{\pi}$ dependence of the long- and intermediate-range parts of the nucleon-nucleon and three-nucleon (3N) potentials is explicit because pion-exchanges are systematically included, in Weinberg power counting. The $m_{\pi}$ dependence of the nuclear interaction energy in PNM and symmetric nuclear matter (SNM) has been studied in the context of addressing how $\langle\bar{q}q\rangle$, the chiral condensate, evolves at finite density Cohen:1991nk ; Kaiser:2007nv ; Plohl:2007dp ; Kruger:2013iza . In these studies, the Feynman-Hellmann theorem was exploited to relate the change in $\langle\bar{q}q\rangle$ to the derivate of the interaction energy with respect to $m_{\pi}^{2}$. Explicitly, one finds that

### v2

@@ II.2 Effective axion potential at finite density
For $\varepsilon=1$, the critical density determined from Eq. (11) is sufficiently large that the expansion in $\sigma_{N}n_{B}/m_{\pi}^{2}f_{\pi}^{2}$ breaks down near the critical density, and it is conceivable that axion condensation does not occur at any density. We will return to this question in Sec. III.2. However, for $\varepsilon<0.4$, Eq. (10) predicts axion condensation at $n_{B}\lesssim n_{\rm sat}$. At these moderate densities where matter can be described reliably using nuclear physics, one can conclude that the axion potential is minimized at $\theta=\pi$. If the nuclear force is significantly modified by axion condensation, $\theta=\pi$ may be favored below the saturation density even for $\varepsilon\geq 0.4$, which we will discuss in more detail in Sec. III and Sec. IV.2.
@@
where $P(\mu_{B},\theta)$ is the pressure, $k_{F}(\theta)=\sqrt{\mu_{B}^{2}-m^{2}_{n}(\theta)}$ and the neutron mass $m_{n}(\theta)$ is given in Eq. (II.1). From Eq. (12) we see that axion contribution to the pressure is negative, and for $\theta=\pi$, $P_{\rm axion}=-V(\theta=\pi)=\varepsilon f_{\pi}^{2}[m_{\pi}^{2}(\theta=\pi)-m_{\pi}^{2}(\theta=0)]$. The importance of this negative pressure was first pointed out in Ref. [63, 37] where it was noted that for white dwarfs, which are stabilized by electron degeneracy pressure, this negative pressure would destabilize a large number of observed white dwarfs in the universe. This consideration leads to the strongest constraint on the axion parameter space for $\varepsilon<2\times 10^{-7}$ [37]. A neutron star, on the other hand, is much denser and balanced by the much larger neutron degeneracy pressure. It is unlikely that this negative pressure can lead to the collapse of a neutron star into a black hole, but it might significantly change the EOS of a neutron star [45].
@@
We begin by analyzing the effects of axions on a dilute gas of neutrons. The left panel of Fig. 1 shows the pressure of a free Fermi gas of neutrons at $\theta=\pi$ and $\mu_{B}=m_{n}(\theta=0)$. The insets show the $\theta$-dependence of the free energy in each of the regions depicted. When the pressure is positive, axion condensation will be favored for dilute neutron matter. In the absence of nuclear interactions, dilute neutron matter with $\theta=0$ is metastable for $0.005\lesssim\varepsilon\lesssim 0.065$ and unstable when $\varepsilon\lesssim 0.005$. For $\varepsilon\gtrsim 0.065$, the first-order transition to the $\theta=\pi$ phase occurs at $\mu_{B}>m_{n}$ and low-density neutron matter with $\theta=0$ is stable. In this case, the phase transition to axion condensed matter will occur at finite density. On the other hand, if $\varepsilon\lesssim 0.065$, the phase transition will occur at $\mu_{B}<m_{n}$, and bulk matter in its ground state will always have $\theta>\theta_{c}$, with $m_{n}(\theta_{c})=\mu_{B}$. In the case where $2\times 10^{-7}<\varepsilon\lesssim 0.005$, normal matter is unstable, and ordinary stars, planets, and white dwarfs are stabilized by Coulomb forces between atoms and nuclei. For $\varepsilon>2\times 10^{-7}$, a system of atoms or nuclei at fixed chemical potential must significantly increase its density to nucleate the $\theta=\pi$ phase, and an electrical barrier must be overcome to bring the nuclei closer together. For smaller $\varepsilon$, the required increase in density is more modest, and the constraint on axions from the stability of white dwarfs, stars, and planets applies.
For a given $\varepsilon$, if $\mu_{B}$ inside the star exceeds the critical chemical potential and it is possible to nucleate the $\theta=\pi$ phase, axions would condense in the inner region, separated by a domain wall with thickness $\Delta r_{\rm DW}\approx 1/m_{a}$ from the exterior $\theta=0$ phase. In the Fermi gas model, for $\varepsilon\gtrsim 0.065$, the neutron star would consist of three regions: a normal outer region with $\theta=0$, a metastable region which could, in principle, exist in the $\theta=0$ phase if the nucleation timescale was long, and inner core characterized by $\theta=\pi$. In the right panel of Fig. 1, we plot pressure as a function of $\mu_{B}$ for $\varepsilon=0.3$. It consists of the three regions just mentioned, and the insets show the $\theta$ dependence of the free energy, which delineates the stable, metastable, and unstable regions.
II.3 Axion field profile interpolating between $\theta=\pi$ and $\theta=0$
@@
where $\mu^{c}_{B}$ is the critical baryon chemical potential at which the transition occurs. This equation, though used here for a dilute gas of nucleons, remains exact when nuclear interaction is included in $\Omega(\theta,\mu^{c}_{B})$. Such an axion profile will develop as long as the object that contains this dilute gas of nucleons is both dense enough (see Eq. (11)) and its radius, $r_{\rm NS}$, large enough such that the bulk of the star can be subsumed within a domain wall of thickness
$\Delta r_{\rm DW}\simeq\frac{1}{|m_{a,\rm eff}(n_{B})|}\leq r_{\rm NS}.$ (14)
@@
We will focus on the condition on the density in Eq. (11) for now and return to the effects of the finite size of stellar objects in Sec. V. For axions with $m_{a}<10^{-11}\,{\rm eV}$, the axion profile can be treated as situated outside the neutron star. In binary systems, this field profile results in axion-mediated forces and axion radiation, which lead to the most stringent constraint from LVK observations of neutron star binaries [39]. On the other hand, when $m_{a}\gg 10^{-11}\,{\rm eV}$, the axion field profile can reside inside the neutron star, separating regions with different densities and pressure. Given the prior constraints on the exceptionally light QCD axion parameter space, we are mostly interested in the region of parameter space where $|m_{a}^{2}(n_{B})|>m_{a}^{2}>1/r_{\rm NS}^{2}$.
@@
For these larger axion masses, the thickness of the region in which $\theta$changes from ±$\pm\pi$ to 0 is much smaller than the radius of a neutron star. On the other hand, constraints on QCD axions from SN cooling imply that the axion mass is below the tens of meV scale [34], i.e. the domain wall thickness is much larger than the relevant microphysical scales: the size of a nucleon, typical separation between nucleons, Debye screening length, and mean free path of neutrons and protons in a neutron star. As a result, we can treat the axion field profile in the parameter space of interest as smooth when considering interactions between the nucleons, while at the same time, studying the properties of the gas of nucleons in domains where $\theta=\pm\pi$ and $\theta=0$ separately, and glue them together with a boundary condition at the location of a thin axion “domain wall.”
@@
Unlike a normal neutron star where density and pressure are continuous, with an axion field profile, both the density and pressure are no longer continuous when adopting the approximation that the domain wall is infinitely thin. In addition to the negative pressure from the condensation of the axion field itself, the domain wall also exerts a force inward due to domain wall tension of order $m_{a}f_{a}^{2}/r_{\rm NS}$. For $m_{a}\gg 1/r_{\rm NS}$, this additional pressure is parametrically smaller in the thin wall limit in the model in Eq. (5). We return to discuss the importance of domain wall pressure in the $\mathbb{Z}_{\mathcal{N}}$ model in Sec. IV.2.
III Neutron and nuclear matter at $\theta=\pi$
@@
In this section, we describe our efforts to model the properties of dense nuclear matter at $\theta=\pm\pi$, taking into account changes to the strength of nuclear interactions. Despite developments in Effective Field Theory (EFT) methods to describe nuclear interactions, the connection between fundamental parameters that appear in the QCD Lagrangian (the quark mass matrix ${\cal M}_{q}$ and $\theta$) and the parameters needed to describe interactions between nucleons remains elusive. This is because the quark mass dependence of the low-energy constants (LECs) in the EFT that describe the strength of interactions at short distances is poorly known. Nevertheless, earlier studies have provided useful insights to address the question of the quark mass dependence of the nucleon-nucleon interaction.
@@
We first briefly summarize earlier work in Refs. [64, 65, 58] which are based on nuclear potentials derived from ChiEFT as they are better suited to address how nuclear interactions are altered by changes to the quark mass matrix. To assess the quark mass dependence of the nuclear interactions at momentum scales of relevance to the large nuclei and neutron-rich matter encountered in neutron stars, we next compare results from ChiEFT and the three scenarios in RMFT which we will use to extract phenomenological results, summarized in Table 1.
[…]
@@
RMFT is based on the assumption that scalar and vector interactions provide the dominant contributions to the interaction energy of baryons in large systems. In this phenomenological model, the interaction between nucleons is mediated by the exchange of scalar and vector mesons. The model and the quark mass dependence of the model parameters are described in Sec. III.3 (see table 1 for a summary). While ChiEFT is a microscopic theory in which potentials are constrained by nucleon-nucleon scattering data and properties of light nuclei, RMFT is a phenomenological model designed to provide a good description of large nuclei. RMFT also provides a convenient extrapolation to high density and permits efficient calculation of heterogeneous crust phases. In RMFT Scenario A, the only modification we make to nuclear physics is to modify the nucleon masses according to Eq. (II.1) since the pion does not explicitly appear in RMFT. In Scenario B, we modify the meson masses with $\theta$ dependence given by calculations in Refs. [56, 57, 58]. In Scenario C, we consider a more extreme modification to the RMFT based on the results around saturation density given by ChiEFT.
[…]
Scenario $K_{\sigma}$ interpretation
A 00 Nuclear forces are not modified
B $0.08$ Corresponds to an $m_{\sigma}$ decrease of $\sim 6\%$ as found in Refs. [56, 57, 58]
C $0.16$ Matches the binding of neutron matter found using ChiEFT. See Figs. 2 & 4.
@@
III.1 Pion mass dependence of two-nucleon observables
By explicitly accounting for the $m_{\pi}$ dependence of the parameters that describe the nucleon-nucleon (NN) interaction in ChiEFT, one can calculate the quark mass dependence of low-energy NN observables such as the NN scattering phase shifts and the deuteron binding energy. In ChiEFT, at leading order (LO) in the momentum expansion, the NN potential is characterized by long-range one-pion exchange (OPE) and short-range NN contact interactions. While the $m_{\pi}$ dependence of the parameters associated with the OPE potential is well understood, the $m_{\pi}$ dependence of the NN contact interaction, which is typically called $D_{2}$, is not well determined and the resulting uncertainty in the scattering length and scattering phase shifts have been studied in earlier work [64, 65].
In Ref. [65], the authors find that when the quark mass is reduced, the binding energy of the deuteron increases, and the spin-singlet and spin-triplet scattering lengths decrease. To quantify the quark mass dependence of an observable $\cal O$, the authors define the logarithmic derivative at the physical point
@@ III.1 Pion mass dependence of two-nucleon observables
since at leading order in ChiPT, $m_{\pi}^{2}\propto m_{q}$. The $m_{q}$ dependence of the ${}^{1}S_{0}$ and ${}^{3}S_{1}$ scattering lengths and the deuteron binding energies are captured by the values of $K_{a1S0}$, $K_{a3S1}$, and $K_{\rm deut}$, and their values obtained in earlier studies are shown in Table 2. Results shown in the first row were obtained using potentials derived from ChiEFT at next-to-leading order (NLO) in Weinberg power counting [65] and NLO results obtained in Ref. [64] using Kaplan-Savage-Wise power counting [66] are shown in the second row. In this case, the uncertainty is dominated by the poorly known LEC, $D_{2}$, which was mentioned earlier. Results obtained more recently in Ref. [58] are shown in the third row; here, the authors employ strategies to go beyond the NLO approach adopted in Refs. [64, 65] and use information from resonance saturation models to constrain the quark mass dependence of the short-range contact operators and incorporate N2LO corrections to $g_{A}$, $f_{\pi}$, and the nucleon mass.
Table 2: Quark mass dependence of the scattering lengths and the deuteron binding energy $K_{\cal O}$, as defined in Eq. (15).
@@
Epelbaum et al. [65] $5\pm 5$ $1.1\pm 0.6$ $-2.8\pm 1.2$
Beane et al. [64] $2.4\pm 3.0$ $3.0\pm 3.5$ $-7\pm 6$
Berengut et al. [58] $2.3^{+1.6}_{-1.5}$ $0.32^{+0.17}_{-0.18}$ $-0.86^{+0.45}_{-0.50}$
With decreasing quark mass, these earlier studies found the ${}^{1}S_{0}$ phase shifts were reduced and the ${}^{3}S_{1}$ phase shifts were enhanced. Thus, in dilute neutron matter where spin singlet interactions dominate, the interaction energy will be less attractive when s-wave interactions dominate. In dilute nuclear matter with an equal number of protons and neutrons (symmetric nuclear matter), where spin triplet interactions dominate, the s-wave interaction energy is larger and more attractive when the quark mass is reduced. However, when the pion becomes lighter and of longer range, higher partial wave contributions are in general enhanced [67]. The attraction in the ${}^{3}P_{0}$ and ${}^{3}P_{2}$ channels is enhanced and the repulsion in the ${}^{1}P_{1}$ channel is reduced [65]. For these reasons, the analysis of the scattering lengths and the deuteron binding energy is insufficient to inform about the behavior of matter at densities of relevance to nuclei and neutron stars where contributions due to higher partial waves and higher-order quark mass-dependent short-distance operators can contribute.
III.2 Axion condensation in Chiral Effective Field Theory
To address how nuclear interactions influence axion condensation, we first make the simple observation that axion condensation is favored when the interaction energy per particle decreases with pion mass. From the discussion in Sec. II, we infer that in this case, condensation would occur at $n_{B}\lesssim 2.6n_{\rm sat}$ for $\varepsilon=1$. In ChiEFT, the $m_{\pi}$ dependence of the long- and intermediate-range parts of the nucleon-nucleon and three-nucleon (3N) potentials is explicit because pion-exchanges are systematically included in Weinberg power counting. The $m_{\pi}$ dependence of the nuclear interaction energy in PNM and symmetric nuclear matter (SNM) has been studied in the context of addressing how $\langle\bar{q}q\rangle$, the chiral condensate, evolves at finite density [68, 69, 70, 71]. In these studies, the Feynman-Hellmann theorem was exploited to relate the change in $\langle\bar{q}q\rangle$ to the derivative of the interaction energy with respect to $m_{\pi}^{2}$. Explicitly, one finds that